5x+3y=30 Find y-intercept and the x-intercept for the equation...
Answer by nerdybill(7003)

5x+3y=30 Find y-intercept and the x-intercept for the equation...
.
For y-intercept, set x=o and solve for y:
5x+3y=30
5(0)+3y=30
0+3y=30
3y=30
y = 10
So, y-intercept is at (0, 10)
.
For x-intercept, set y=o and solve for x:
5x+3y=30
5x+3(0)=30
5x+0=30
5x=30
x = 6
So, x-intercept is at (6,0)